# Incompressible Limit of Compressible Ideal MHD Flows inside a Perfectly Conducting Wall

### v4

@@ 4 Incompressible limit
for $\delta>0$ because $\|\varepsilon\partial_{t,x}p\|_{3}$ and $\|u\|_{4}$ are uniformly bounded in $[0,T]$. Now, $(u^{0},B^{0},S^{0})\in L^{\infty}([0,T];H^{4}(\Omega))$ solves the incompressible MHD equations together with a transport equation
@@
$\begin{cases}\varrho(\partial_{t}u^{0}+u^{0}\cdot\nabla u^{0})-B^{0}\cdot\nabla B^{0}+\nabla(\pi+\frac{1}{2}|B^{0}|^{2})=0&~{}~{}~{}\text{in}~{}[0,T]\times\Omega,\\ \partial_{t}B^{0}+u^{0}\cdot\nabla B^{0}-B^{0}\cdot\nabla u^{0}=0&~{}~{}~{}\text{in}~{}[0,T]\times\Omega,\\ \nabla\cdot u^{0}=\nabla\cdot B^{0}=0&~{}~{}~{}\text{in}~{}[0,T]\times\Omega,\\ \partial_{t}S^{0}+u^{0}\cdot\nabla S^{0}=0&~{}~{}~{}\text{in}~{}[0,T]\times\Omega,\\ u_{3}^{0}=B_{3}^{0}=0&~{}~{}~{}\text{on}~{}[0,T]\times\Sigma,\\ (u^{0},B^{0},S^{0})|_{t=0}=(u_{0}^{0},B_{0}^{0},S_{0}^{0}),\end{cases}$ (4.3)

### v6

@@ 4 Incompressible limit
for $\delta>0$ because $\|\varepsilon\partial_{t,x}p\|_{3}$ and $\|u\|_{4}$ are uniformly bounded in $[0,T]$.
Now, following the argument in Klainerman-Majda [14, Theorem 1], there exists a function $\pi$ satisfying $\nabla\pi\in C([0,T];H^{3}(\Omega))$ such that
  1. 1.
    The following convergence result holds
    $\rho^{-1}\nabla p\to\varrho^{-1}\nabla\pi\quad\text{weakly-* in }L^{\infty}([0,T];H^{3}(\Omega))\text{ and strongly in }C([0,T];H^{3-\delta}(\Omega)).$
  2. 2.
@@
    $(u^{0},B^{0},S^{0})\in L^{\infty}([0,T];H^{4}(\Omega))$ solves the incompressible MHD equations together with a transport equation
@@
    $\begin{cases}\varrho(\partial_{t}u^{0}+u^{0}\cdot\nabla u^{0})-B^{0}\cdot\nabla B^{0}+\nabla(\pi+\frac{1}{2}|B^{0}|^{2})=0&~{}~{}~{}\text{in}~{}[0,T]\times\Omega,\\ \partial_{t}B^{0}+u^{0}\cdot\nabla B^{0}-B^{0}\cdot\nabla u^{0}=0&~{}~{}~{}\text{in}~{}[0,T]\times\Omega,\\ \nabla\cdot u^{0}=\nabla\cdot B^{0}=0&~{}~{}~{}\text{in}~{}[0,T]\times\Omega,\\ \partial_{t}S^{0}+u^{0}\cdot\nabla S^{0}=0&~{}~{}~{}\text{in}~{}[0,T]\times\Omega,\\ u_{3}^{0}=B_{3}^{0}=0&~{}~{}~{}\text{on}~{}[0,T]\times\Sigma,\\ (u^{0},B^{0},S^{0})|_{t=0}=(u_{0}^{0},B_{0}^{0},S_{0}^{0}).\end{cases}$ (4.3)
    Here $\varrho$ satisfies $\partial_{t}\varrho+u^{0}\cdot\nabla\varrho=0,$ with initial data $\varrho_{0}:=\rho(0,S_{0}^{0})$.